How many lithium ions are present in 25.0 mL of 0.300 M Li2CO3 solution?
a. 9.03 x 1022 ions
b. 4.52 x 1021 ions
c. 0.0150 ions
d. 9.03 x 1021 ions
1
Expert's answer
2020-03-26T03:02:52-0400

C(Li2CO3) = 0.3 M

V(solution) = 25 mL = 0.025 L

n(Li2CO3) = C*V = 0.0075 mol

n(Li+) = 2n(Li2CO3) = 0.0075*2 = 0.015 mol (because Li2CO3 has 2 lithium ions in its formula)

N(Li+) = n(Li+)*NA = 0.015*6.02*1023 = 9.03*1021

correct answer is d